OpenStudy (anonymous):

(cos pi/4 - tan pi/4)^2

OpenStudy (anonymous):

Find the value of : \[\cos(\frac{\pi}{4}) = ??\]

OpenStudy (anonymous):

@yasmin95 u know the value of cos and tan

OpenStudy (anonymous):

Then find : \[\tan(\frac{\pi}{4}) = ??\]

OpenStudy (anonymous):

(1/square root 2 - 1)^2

OpenStudy (anonymous):

Then use : \[(a-b)^2 = a^2 + b^2 - 2ab\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Yes good.. Go ahead now..

OpenStudy (anonymous):

\[(\frac{1}{\sqrt{2}} - 1)^2 \implies \frac{(1 - \sqrt{2})^2}{(\sqrt{2})^2}\]

OpenStudy (anonymous):

In the numerator, use the above formula..

OpenStudy (anonymous):

how did u get the above step ?

OpenStudy (anonymous):

This is by using LCM or LCD concept: \[\frac{1}{\sqrt{2}}- 1 \implies \frac{1}{\sqrt{2}}- \frac{\sqrt{2}}{\sqrt{2}} \implies \frac{1 - \sqrt{2}}{\sqrt{2}}\]

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

i solved it thanks !!

OpenStudy (anonymous):

More explanation??

OpenStudy (anonymous):

Okay.. You are welcome dear..